# SSR: Alignment-Aware Modality Connector for Speech Language Models

### v1

@@ 3.2 Speech-text Aligners
For both CTC and UnitY2 aligners, we extract segmentations from the alignments and then apply selection-based compression. We also experimented with Continuous Integrate-and-Fire (Dong & Xu, 2020, CIF) as the connector, which is designed to learn segmentation and perform compression simultaneously. Instead of relying on a fixed, pre-computed segmentation, CIF dynamically segments and aggregates speech features by scoring each feature and computing a weighted average. For more details, we refer readers to the original paper (Dong & Xu, 2020).
3.3 Training Method
With SSR-Connector, we convert speech into representations with the same granularity as their transcription tokens. This allows us to fine-tune the SpeechLM directly using the next-token prediction objective, where the input is the compressed representation $\bm{z}$ and the target is the transcription $\bm{y}$. This approach is possible because our feature $\bm{z}$ and text token $\bm{y}$ share the same length $m$. However, our preliminary studies showed that directly fine-tuning with the next-token prediction objective leads to catastrophic forgetting, undermining the pre-trained LLM’s abilities. Therefore, we propose a two-stage training pipeline consisting of a distillation stage and a fine-tuning stage, as shown in Fig. 3.
In Stage 1, we pre-train SSR-Connector by distilling the LLM’s text embeddings to align the connector’s representations with the LLM’s embedding space. Formally, given aligned speech-text data, we compute the text embeddings $\bm{h}=f(\bm{y};\theta_{\text{emb}})$, where $\bm{y}$ is the transcription token sequence, $\theta_{\text{emb}}$ is the embedding table, and $f$ maps tokens $\bm{y}$ to their embeddings. Following our connector design in § 3.1, we obtain the compressed speech representations $\bm{z}$. For distillation, we use a combination of cosine similarity loss $\mathcal{L}_{\text{cos}}$ and mean squared error (MSE) loss $\mathcal{L}_{\text{MSE}}$
$\mathcal{L}=\lambda\mathcal{L}{\text{cos}}+\mathcal{L}_{\text{MSE}}=\frac{1}{m}\sum_{i=1}^{m}\left[\lambda\left(1-\frac{\mathbf{z}_{i}^{\top}\mathbf{h}_{i}}{|\mathbf{z}_{i}|\cdot|\mathbf{h}_{i}|}\right)+\left|\mathbf{z}_{i}-\mathbf{h}_{i}\right|^{2}\right],$ (4)
We offer detailed descriptions of different aligners and demonstrate the performance of SpeechLM after distillation training in § 4. In § 5, we present results after fine-tuning SpeechLM and compare various fine-tuning strategies to identify the method that minimizes catastrophic forgetting.
4 Stage 1: Alignment-aware Connector Distillation
4.1 Datasets
For distillation training, we use the aligned speech-to-text dataset MLS (Pratap et al., 2020), specifically the English portion, which consists of about 50,000 hours of speech. To evaluate our SpeechLMs, we employ several datasets as shown in Table 1. To assess the model’s spoken language understanding (SLU) capabilities, we follow Nguyen et al. (2024) and use sWUGGY, sBLIMP, and the StoryCloze dataset. sWUGGY and sBLIMP are detailed in (Nguyen et al., 2020). Briefly, sWUGGY evaluates whether a model can discriminate between real spoken words and non-words (e.g., “brick” vs. “blick”), while sBLIMP assesses if the model can distinguish between a grammatically correct spoken sentence and its ungrammatical variant (e.g., “cats are lazy” vs. “cats is lazy”). We evaluate our SpeechLMs on both text ($T$) and speech ($S$) versions of sWUGGY and sBLIMP. The StoryCloze dataset measures whether the model can identify the plausible ending between two sentences given the beginning of a short story, which typically requires high-level semantic understanding and common sense (Mostafazadeh et al., 2017). Besides spoken and text versions of StoryCloze, following Nguyen et al. (2024), we use a speech-text version ($S\rightarrow T$), where the beginning of the story is synthesized into speech and the two ending sentences are kept in text format. This version requires the model to have cross-modal understanding to infer the sensible story ending.
MMLU (Hendrycks et al., 2021) is widely used to assess LLMs’ knowledge comprehension, understanding, and reasoning abilities, and we use it to measure the extent of forgetting during cross-modal fine-tuning. Since MMLU is a diverse and high-quality evaluation dataset for LLMs, we craft a variant, Speech-MMLU, to assess our SpeechLM’s cross-modal understanding. Specifically, we utilized AudioBox (Vyas et al., 2023), a high-quality text-to-speech synthesizer, to convert the question portion of each choice task into speech while keeping the multiple-choice answers in text format. We selected a subset of MMLU to construct our Speech-MMLU dataset, as some domains’ questions are not suitable for synthesis (e.g., the algebra subset contains many mathematical notations that are not synthesized properly). sWUGGY, sBLIMP, StoryCloze, and Speech-MMLU are all categorized as ”Choice Task”, meaning several choices are presented to the SpeechLM (Speech-MMLU has four choices while the other task has only two choices). For each task, we compute accuracy using groundtruth choice and the highest likelihood choice predicted by the SpeechLM.
Lastly, we also evaluate our SpeechLM’s ASR performance using the Librispeech clean/other datasets. We evaluate ASR in a prompt-based fashion with zero-shot and five-shot setting. More details about our evaluation (e.g., prompts for ASR, Speech-MMLU construction, etc.,) can be found in Appendix.
Eval Dataset Type Eval Metric Eval Format
sWUGGY (Nguyen et al., 2020) Choice Task Accuracy $S,T$
sBLIMP (Nguyen et al., 2020) Choice Task Accuracy $S,T$
MMLU (Hendrycks et al., 2021) Choice Task Accuracy $T$
Speech-MMLU (Ours) Choice Task Accuracy $S\rightarrow T$
4.2 Model Setup
We instantiate our LLM using the pre-trained LLaMa3 model (Touvron et al., 2023) and employ DinoSR (Liu et al., 2023a) as our pre-trained speech feature extractor. Our speech connector includes a linear layer that maps DinoSR’s extracted representations ($D=768$) to the LLM’s embedding space dimension ($H=4096$). We then utilize a 4-layer Transformer Decoder to transform and compress the speech representations based on alignments, as described in § 3.1. The compressed representations $\bm{z}$ and the embeddings of text tokens $\bm{h}$ are used to compute the distillation loss for updating the connector’s parameters. We train our connector for 400,000 steps with a learning rate of $1\times 10^{-5}$, using dynamic batching with a maximum of 4,096 tokens per device. We employ distributed data parallelism (DDP) with 32 A100 GPUs.
To extract alignments, we experimented with different aligners listed in § 3.2. For the UnitY aligner222 Publicly available at https://github.com/facebookresearch/seamless_communication/blob/main/docs/m4t/unity2_aligner_README.md, we used it off-the-shelf to construct alignment indices. Since the UnitY2 aligner provides alignments based on character-level tokens, we merge the durations into subword level to ensure that the compressed representations and text embeddings have the same granularity. For CTC-based aligners, we trained them using a 4-layer Transformer Decoder followed by a linear projection. In the character-level variant (Char-CTC), we deduplicate the sequence to obtain character-level durations and then merge them into subword-level durations to segment the speech features, similar to the UnitY2 aligner. In the subword-level variant (Sub-CTC), we directly use CTC’s blank token to segment the speech input.
4.3 Aligner Performance Comparison
@@
Table 2: Performance comparison (with Cosine Similarity, MSE, and 0/5-shot ASR WER) between different aligners used for Stage 1 training, evaluated on Librispeech clean test set.
To compare the quality of different aligners, we trained several SSR-Connector based on different aligners via distillation. We evaluated the aligners using the Librispeech clean test set by computing the Cosine Similarity (Cos(%)) and Mean Squared Error (MSE) between the compressed representations and text embeddings. Additionally, we performed zero-shot and five-shot ASR with the learned connector. Note that we never explicitly trained the model for ASR tasks, and the base LLM remained frozen during Stage 1 training. Therefore, the model achieves low word error rates (WER) only when the distilled speech representations closely resemble the text embeddings. As shown in Table 2, the UnitY2 aligner brings the speech representations close to their corresponding text embeddings, achieving very low WER in both zero-shot and five-shot ASR settings. Among textless aligners, we found that Char-CTC performs the best, likely because it has a much smaller vocabulary compared to Sub-CTC, making it easier to learn. Lastly, CIF resulted in suboptimal performance, possibly due to its less accurate alignment, as its segmentation is predicted by accumulating scores without exploiting the monotonicity between speech and text.
To visualize the effect of distillation, we present t-SNE plots of the adapted speech representations and text embeddings in Fig. 4, categorizing them into high and low similarity groups based on the cosine similarity between Char-CTC representations and text embeddings. We observe that longer subwords tend to exhibit higher similarity, likely because their long segments make it easier for the connector to convert speech representations into corresponding text embeddings. Furthermore, longer subwords possess more coherent semantics compared to shorter tokens like ‘wy’ or ‘ia’.
@@
Table 3: Alignment quality of aligners.
@@
Given that UnitY2 and Char-CTC performs the best, we also follow Huang et al. (2024) to measure their word boundary error (WBE) and word average duration (WDUR) using the TIMIT (Garofolo et al., 1993) data. Though the aligner quality can be further improved with other methods such as CTC + Label Prior (Huang et al., 2024), MMS (Pratap et al., 2023), or MFA (McAuliffe et al., 2017), Char-CTC and UnitY2 still achieve good quality and we choose them out of simplicity and general availability (unlike ”CTC+Label Prior”, for example, which requires customization with library like k2333https://github.com/k2-fsa/k2).
4.4 Experimental Results
In the previous section (§ 3.2), we compared different aligners and found that UnitY2 and Char-CTC performed the best. Consequently, we evaluate SpeechLM using these two aligners. First, we assess the model on Spoken Language Understanding (SLU) tasks and the MMLU benchmark (Hendrycks et al., 2021). We compare our model against several generative speech systems, all of which utilize Transformer-Decoder models trained on speech units. These methods vary in training approaches (pretrained from scratch or fine-tuned), types of speech units, and the size of training data.
Briefly, GSLM (Lakhotia et al., 2021) trains on speech units like HuBERT, TWIST (Hassid et al., 2024b) is a textually pretrained speech model based on Llama-13B (Touvron et al., 2023), and AudioLM (Borsos et al., 2023) employs a cascade system with a semantic sequence model alongside coarse- and fine-acoustic models. These models focus solely on speech without capabilities for text understanding or generation. More recently, SpiritLM (Nguyen et al., 2024) and VoxtLM (Maiti et al., 2024) have adopted multi-task training objectives that incorporate text-only, speech-only, and speech-text token sequences to fuse the speech modality into pre-trained LLMs effectively. Since the original SpiritLM is fine-tuned based on LLaMa2, we follow the same recipe to fine-tune the LLaMa3-based SpiritLM ourself for a fair comparison on text relevant metrics like MMLU.
Model Type sWUGGY sBLIMP Storycloze MMLU
[…]
   Previous Work
Voxtlm (Maiti et al., 2024) 80.3 66.1 74.2 57.1
TWIST (Hassid et al., 2024b) $\emptyset$ 74.5 $\emptyset$ 59.2 $\emptyset$ 55.4 $\emptyset$ $\emptyset$
SpiritLM (Nguyen et al., 2024) 80.3 69 73.3 58.3 79.4 61 64.6 36.9
SpiritLM (LLaMa3) 77.6 73.5 74.5 56.3 75.1 61.1 61.6 53.5
  SSR-Connector
UnitY2 + Blockwise-mask 81 71.5 74.5 73.1 80.9 71.8 75 65.3
UnitY2 81 71.2 74.5 72.4 80.9 69.3 74.8 65.3
Char-CTC 81 56.4 74.5 67.3 80.9 62.2 74.3 65.3
Char-CTC (Unit-based) 81 54.1 74.5 61.8 80.9 59.2 72.5 65.3
   Cascade System
ASR (Whipser) + LLaMa2 84.1 79.2 72.8 71.6 81.9 75.7 75.7 46.2
Table 4: Model performance on spoken language understanding and MMLU. : Results taken from Nguyen et al. (2024).: Results taken from Défossez et al. (2024). : Our implementation of SpiritLM based on LLaMa3 checkpoint. We fill with $\emptyset$ the task and modality that are not supported by the reported system, and with    the scores that are not publicly available. We bold the best result and highlight the second-best system with the blue color box (excluding the cascaded system).
Spoken Language Understanding Performance
Beyond UnitY2 and Char-CTC, we introduce two additional systems for ablation. The UnitY2 + Blockwise-mask system achieves the highest performance by applying a blockwise attention mask to further constrain the Transformer-Decoder (described in § 3.1). However, due to its marginal improvement over UnitY2 and increased computational cost, we decide to simplify the design and remove the blockwise-attention masks. The Char-CTC (Unit-based) system differs by utilizing discrete speech units instead of raw waveform features processed by the DinoSR (Liu et al., 2023a) encoder. These units are extracted via K-Means clustering on DinoSR representations, which leads to some information loss during discretization and reconstruction, resulting in lower performance compared to Char-CTC. Nonetheless, Char-CTC (Unit-based) demonstrates that our alignment-aware connector design is compatible with both continuous waveforms and discrete speech units.
Speech-MMLU and Prompt-based ASR Performance
In addition to SLU tasks, we evaluate our systems on the Speech-MMLU benchmark, which assesses cross-modal understanding and is more challenging than previous SLU tasks. We also conduct prompt-based ASR evaluations to assess the quality of the adapted features. As shown in Table 5, our systems greatly outperform the previous SpeechLM (SpiritLM), achieving a +20 accuracy improvement on the Speech-MMLU dataset444 We report micro-average across 22 domains and the detailed breakdown is available in Appendix C.. These results indicate that SpeechLM based on SSR-Connector possesses enhanced cross-modal abilities that enable it to comprehend spoken questions and reason through multiple-choice options to select correct answers. Similarly, our systems achieve much lower WERs on both the Librispeech clean and other test sets compared to SpiritLM. Notably, neither SpiritLM nor our system were trained on ASR tasks, so the model relies solely on in-context learning to generate transcriptions. Even our weakest system (Char-CTC (Unit-based)) can outperform SpiritLM ’s 10-shot result.
0-shot 5-shot 0-shot 5-shot 0-shot 5-shot
SpiritLM (Nguyen et al., 2024) N/A N/A N/A 21.9 N/A 29.2
SpiritLM (LLaMa3) 40.5 42.75 N/A 21.0 N/A 28.5
  SSR-Connector
UnitY2 + Blockwise-mask 65.0 69.5 5.0 2.6 8.1 6.8
UnitY2 64.2 68.6 5.6 4.0 12.1 10.6
Char-CTC 61.7 66.5 9.7 6.5 20.2 14.9
Char-CTC (Unit-based) 57.4 62.3 12.6 8.8 25.6 18.6
Table 5: Comparison of Speech-MMLU and ASR performance. Speech-MMLU results are micro-averages across all domains. : For SpiritLM, We report WER using 10-shot prompting, following Nguyen et al. (2024). N/A: We did not evaluate SpiritLM in those settings.
5 Stage 2: Speech Language Model Fine-tuning
Model Type sWUGGY sBLIMP Storycloze MMLU
[…]
Char-CTC 81 56.4 74.5 67.3 80.9 62.2 74.3 65.3
+ Vanilla Fine-tuning 82.5 56.6 75.8 68.8 75.2 62.8 71 57.4
+ LoRA Fine-tuning 82.4 56.5 75.8 68.7 76.3 62.6 71.5 58.2
+ Multitask Fine-tuning 82.9 56.7 75.9 68.9 81 63.4 73.1 63.1
Table 6: Comparison of different Stage 2 fine-tuning methods (reported after fine-tuned for 5k updates). Multitask fine-tuning obtains the best improvement on SLU tasks while achieving least catastrophic forgetting. We bold the best performance and use blue color box for the second-best result.
5.1 Mitigate Catastrophic Forgetting
Model and Dataset Setup
  • Multitask Fine-tuning: To preserve the LLM’s pre-trained text capabilities, we also fine-tune SpeechLM on text-only data using the standard Negative Log-Likelihood (NLL) loss. The dataloader is configured to sample from both speech-text and text-only datasets with equal probability. We continue using the MLS dataset for speech-text training and utilize a subset of the LLaMa2 training datasets (Touvron et al., 2023) for text-only training.
5.2 Comparison of Fine-tuning Methods
Since model performance does not further improve with additional training steps (as shown in Fig. 5), we utilize the checkpoint trained for 5,000 updates to compare with baseline models. The results are presented in Table 6 and Table 7. Note that even with only 5,000 updates, the model has observed all speech-text data due to our large effective batch size. Across SLU, MMLU, and ASR tasks, the fine-tuned SpeechLM outperforms baseline methods on tasks primarily relying on speech-only information (sWUGGY, sBLIMP, ASR), with multitask fine-tuning achieving the best performance among all fine-tuning methods. However, we also observe a decline in performance on $S\rightarrow T$ tasks such as Speech-MMLU and StoryCloze, indicating that there is still unavoidable degradation of text capabilities which adversely affects SpeechLM’s cross-modal performance.
[…]
0-shot 5-shot 0-shot 5-shot 0-shot 5-shot
SpiritLM (LLaMa3) 40.5 42.75 N/A 21.0 N/A 28.5
Char-CTC 61.7 66.5 9.7 6.5 20.2 14.9
+ Multitask Fine-tuning 48.1 56.3 N/A 5.7 N/A 13.1
Table 7: Speech-MMLU and ASR performance of different models. : For SpiritLM, We report WER using 10-shot prompting for ASR, following Nguyen et al. (2024). N/A: The 0-shot generation of our fine-tuned SpeechLM tends to have hallucinations (keep generating after completing the transcription) so we only report its 5-shot performance.
6 Conclusion
We propose SSR-Connector to inject speech representation into pre-trained LLMs. Through explicitly leveraging speech-text alignment, our connector compresses long and sparse speech information to the same granularity as text tokens. To mitigate catastrophic forgetting, we propose a two-stage training pipeline for modality fusion. Compared to previous baselines, our SpeechLM achieves much better speech understanding ability while retaining its pre-trained text ability.
Supplementary Material
Appendix A Datasets
@@
As described in § 4.1, we employ sWUGGY, sBLIMP, StoryCloze, MMLU, Speech-MMLU and Librispeech datasets to assess model performance. In this section, we provide more examples for each evaluation set. sWUGGY and sBLIMP are simple tasks where two choices can be directly compared. As shown in Table 8, sWUGGY provides two choices that requires models to discriminate real words from non-words. sBLIMP assesses whether model can distinguish between a grammatically correct sentence and its ungramatical variant. MMLU and StoryCloze, on the other hand, have a prefix and choices. The StoryCloze dataset measures whether the model can identify the logical ending between two sentences given the beginning of a short story. Since StoryCloze has a shared prefix, we can synthesize only the prefix part into speech and keep choices in text format, resulting in our $S\rightarrow T$ format evaluation that assess model’s cross-modal understanding. Similarly, for MMLU, we also synthesize its prefix (the question portion) into speech and keep the choices in text format, resulting in our Speech-MMLU dataset. Since some topics have bad audio synthesis quality (e.g., the algebra subset contains many mathematical notations), we only keep 22 topics in our test suite (Table 9).
@@
MMLU During the period when life is believed to have begun, the atmosphere on primitive Earth contained abundant amounts of all the following gases except {”A”: ”oxygen”, ”B”: ”hydrogen”, ”C”: ”ammonia”, ”D”: ”methane”}
@@
Appendix B Evaluation Metric and Prompt

### v2

@@ A.1 Aligner Description
For both CTC and UnitY2 aligners, we extract segmentations from the alignments and then apply selection-based compression Tan et al. (2024). We also experimented with Continuous Integrate-and-Fire (Dong and Xu, 2020, CIF) as the connector, which is designed to learn segmentation and perform compression simultaneously. Instead of relying on a fixed, pre-computed segmentation, CIF dynamically segments and aggregates speech features by scoring each feature and computing a weighted average. For more details, we refer readers to the paper (Dong and Xu, 2020).
A.2 Aligner Performance Comparison
To compare the quality of different aligners, we trained several SSR-Connector based on different aligners via distillation. We evaluated the aligners using the Librispeech clean test set by computing the Cosine Similarity (Cos(%)) and Mean Squared Error (MSE) between the compressed representations and text embeddings. Additionally, we performed zero-shot and five-shot ASR with the learned connector. Note that we never explicitly trained the model for ASR tasks, and the base LLM remained frozen during Stage 1 training. Therefore, the model achieves low word error rates (WER) only when the distilled speech representations closely resemble the text embeddings. As shown in Table 6, the UnitY2 aligner brings the speech representations close to their corresponding text embeddings, achieving very low WER in both zero-shot and five-shot ASR settings. Among textless aligners, we found that Char-CTC performs the best, likely because it has a much smaller vocabulary compared to Sub-CTC, making it easier to learn. Lastly, CIF resulted in suboptimal performance, due to its less accurate alignment, as its segmentation is predicted by accumulating scores without exploiting the monotonicity between speech and text.
@@ A.2 Aligner Performance Comparison
Table 6: Performance comparison (with Cosine Similarity, MSE, and 0/5-shot ASR WER) between different aligners used for Stage 1 training, evaluated on Librispeech.
To visualize the effect of distillation, we present t-SNE plots of the adapted speech representations and text embeddings in Fig. 5, categorizing them into high and low similarity groups based on the cosine similarity between Char-CTC representations and text embeddings. We observe that longer subwords tend to exhibit higher similarity, likely because their long segments make it easier for the connector to convert speech representations into corresponding text embeddings. Furthermore, longer subwords possess more coherent semantics compared to shorter tokens. like ‘wy’ or ‘ia’.
@@
Table 7: Alignment quality comparison.
@@
Given that UnitY2 and Char-CTC performs the best, we also follow Huang et al. (2024) to measure their word boundary error (WBE) and word average duration (WDUR) using the TIMIT (Garofolo et al., 1993) data. Though the aligner quality can be further improved with other methods such as CTC + Label Prior (Huang et al., 2024), MMS (Pratap et al., 2023), or MFA (McAuliffe et al., 2017), Char-CTC and UnitY2 still achieve good quality and we choose them out of simplicity and general availability (unlike "CTC+Label Prior", for example, which requires customization with library like k2333https://github.com/k2-fsa/k2).
Appendix B Beyond Semantics: Speech Style Recognition with In-context Learning
To explore the non-semantic capabilities of our SpeechLM, particularly its ability to retain and utilize paralinguistic information, we conducted additional experiments focusing on speech style recognition through in-context learning. Specifically, we investigated whether the SSR-Connector-based SpeechLM (based on the UnitY2 aligner), can differentiate between various speech styles without explicit training on paralinguistic cues.
We utilized the Expresso dataset Nguyen et al. (2023), which comprises speeches delivered in distinct styles such as happy, sad, whispering, and laughing. Two primary tasks were designed to assess the model’s performance:
  1. 1.
    Whisper vs. Laugh: The model was tasked with identifying whether a given speech was whispered or laughed. The prompt provided to the model was:
    "You are given speeches from two styles. Your task is to judge if the speech is a whisper or laugh. Here are some example speeches: [Speech]: {speech} [Style]: {whisper/laugh}…"
  2. 2.
    Happy vs. Sad: The model was asked to determine if the speech was delivered happily or sadly. The prompt used was:
    "Listen to the following speech and judge if the speaker is happy or sad. Here are some examples: [Speech]: {speech} [Emotion]: {happy/sad}…"
For each task, we evaluated the model’s performance using varying numbers of in-context examples: 0-shot, 1-shot, 5-shot, and 10-shot. The results, averaged over 10 runs, are presented in Table 8. Additionally, we benchmarked a cascaded system comprising Whisper and Llama3 for comparison (this cascaded baseline does no preserve non-semantic information and can only infer the speech style through transcripted content).
Task Model 0-shot 1-shot 5-shot 10-shot
Whisper vs. Laugh Cascaded System 51.6 52.1 52.2 54.7
Ours 49.6 62.4 64.0 75.9
Happy vs. Sad Cascaded System 50.0 51.4 51.8 51.0
Ours 51.6 52.1 52.2 54.7
Table 8: Accuracy of Speech Style Recognition Tasks with In-context Learning
The results indicate that with zero-shot prompting, our model generates predictions close to random chance, as it has not been trained to utilize paralinguistic information. However, with the introduction of a few-shot learning approach, the model significantly improves its ability to distinguish between whispering and laughing speech, achieving up to 75.9% accuracy with 10-shot examples. This suggests that the model’s representations inherently contain paralinguistic information that can be harnessed through in-context learning. For the Happy vs. Sad task, the improvement is modest, peaking at 54.7% accuracy with 10-shot examples. This lesser performance compared to the Whisper vs. Laugh task may be attributed to the subtler differences in emotional expression compared to the more pronounced style differences between whispering and laughing.
Overall, these findings demonstrate that our SpeechLM can effectively leverage in-context learning to recognize different speech styles, thereby highlighting the presence of paralinguistic information within the model’s representations. This capability complements existing methods that incorporate paralinguistic information, such as the use of expressive tokens in SpiritLM Nguyen et al. (2024) or emotion-relevant instruction tuning in SALMONN Tang et al. (2024).
Appendix C Datasets
Eval Dataset Type Eval Metric Eval Modality
sWUGGY (Nguyen et al., 2020) Choice Task Accuracy $S,T$
sBLIMP (Nguyen et al., 2020) Choice Task Accuracy $S,T$
MMLU (Hendrycks et al., 2021) Choice Task Accuracy $T$
Speech-MMLU (Ours) Choice Task Accuracy $S\rightarrow T$
@@ Appendix C Datasets
As described in § 4.1, we employ sWUGGY, sBLIMP, StoryCloze, MMLU, Speech-MMLU and Librispeech datasets to assess model performance. In this section, we provide more examples for each evaluation set. sWUGGY and sBLIMP are simple tasks where two choices can be directly compared. As shown in Table 10, sWUGGY provides two choices that require models to discriminate real words from non-words. sBLIMP assesses whether the model can distinguish between a grammatically correct sentence and its ungrammatical variant.
MMLU and StoryCloze, on the other hand, have a prefix and choices. The StoryCloze dataset measures whether the model can identify the logical ending between two sentences given at the beginning of a short story. Since StoryCloze has a shared prefix, we can synthesize only the prefix part into speech and keep choices in text format, resulting in our $S\rightarrow T$ format evaluation that assess the model’s cross-modal understanding. Similarly, for MMLU, we also synthesize its prefix (the question portion) into speech and keep the choices in text format, resulting in our Speech-MMLU dataset. Since some topics have bad audio synthesis quality (e.g., the algebra subset contains many mathematical notations), we only keep 22 topics in our test suite (as shown in the “Topic” column of Table 11).
@@
MMLU During the period when life is believed to have begun, the atmosphere on primitive Earth contained abundant amounts of all the following gases except {"A": "oxygen", "B": "hydrogen", "C": "ammonia", "D": "methane"}
@@
Appendix D Evaluation Metric and Prompt